It takes Kay 20 minutes to get to work traveling 45 mph. Two minutes after she left home this morning, her husband, Dan, started out with her briefcase, which she had forgotten. If Dan arrived at Kay's office just as she did, how fast did he drive?
Answer by Alan3354(69443)

It takes Kay 20 minutes to get to work traveling 45 mph.
--------
20 mins = 1/3 hour
d = r*t
---------
Two minutes after she left home this morning, her husband, Dan, started out with her briefcase, which she had forgotten. If Dan arrived at Kay's office just as she did, how fast did he drive?
----
d = r*t
t = 18 minutes = 3/10 hours
r = d/t